Suppose that

𝑇1 =(𝑋1 βˆ’πœ‡)2

𝑇2 =1

2[(𝑋1 βˆ’πœ‡)2 +(𝑋2 βˆ’πœ‡)2]

And

𝑇3 =1

2(𝑋1 βˆ’π‘‹2)2

Are estimators for 𝜎2.

i) Show whether or not 𝑇3 is an unbiased estimator of 𝜎2. (4)

ii) Which estimator is the most efficient between 𝑇1 and 𝑇2? (6)


1
Expert's answer
2021-10-20T17:48:05-0400

i)

"E(\sigma^2)=\frac{n-1}{n}\sigma^2"


"E(T_3)=E(\frac{(X_1-X_2)^2}{2})=\frac{1}{2}(2(\mu^2+\sigma^2)-2\mu^2)=\sigma^2/2"


Since "E(\sigma^2)\neq E(T_3)" , T3 is a biased estimator of 𝜎2.


ii)

"E(T_1)=" "E(T_1)=E()=""E(T_1)=(n-1)\sigma^2"


"E(T_2)=\frac{1}{2}\cdot2(n-1)\sigma^2=(n-1)\sigma^2"


So, 𝑇1 and 𝑇2 are same efficient estimators.